OpenStudy (anonymous):

is e^(-u) equal to 1/e^(u)?

hartnn (hartnn):

yes. x^(-m) = 1/ x^m

OpenStudy (anonymous):

Hey, Oh My God, you are right dear..

OpenStudy (anonymous):

it is also applicable to e?

hartnn (hartnn):

yes, to any variable/number.

OpenStudy (anonymous):

You can think of it as too : \[\frac{1}{e^{-u}} = e^u\]

OpenStudy (anonymous):

ok thanks a lot! :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

And I am not typing... @hartnn